# SCAWaveNet: A Spatial-Channel Attention-based Network for Global Significant Wave Height Retrieval

### v1

@@ V-C NDBC Buoy Data Evaluation
The Buoy data serve as a critical source of ocean observations. Unlike model-generated data, buoy measurements are continuously collected via onboard sensors and instruments, offering higher accuracy and better representation of real-world marine conditions. Therefore, buoy-derived SWH measurements are widely regarded as a reliable reference for ocean state assessment. It is necessary to evaluate model’s generalization and robustness with buoy data. In this study, NDBC buoy data are utilized for model evaluation. Specifically, 42 NDBC buoys were selected, as illustrated in Fig. 10, where red dots mark their locations, predominantly distributed along the eastern and western coasts of America and near the Hawaiian Islands. These data align with the temporal range of the test set. Data from each channel are individually matched based on close temporal and spatial proximity, and then they are merged into four-channel samples, yielding a total of 4,438 four-channel data samples.
@@
Figure 11: Density scatter plots on CYGNSS-Buoy data: (a) CNN-ConvLSTM, (b) ViT-Wave, (c) WaveTransNet, (d) $\text{MLP}_{f}$-CD, (e) $\text{CNN}_{f}$-CD, (f) SCAWaveNet-CI, (g) SCAWaveNet-CD, (h) SCAWaveNet-CD-wind.
TABLE V: Average Performance comparison of models on buoy data
@@
Table V presents the evaluation results of all models on the buoy dataset. As a global reanalysis dataset, ERA5 exhibits high consistency with the NDBC buoy data, with an average CC of 0.954 across four channels. This result confirms the reliability of using ERA5 as a reference for model evaluation. Among the evaluated results, SCAWaveNet-CD achieves the best performance on buoy data, with an average RMSE of 0.432 m and a CC of 0.754. SCAWaveNet-CI closely follows, with an RMSE of 0.447 m and a CC of 0.742. These findings indicate that the SCAWaveNet architecture effectively captures meaningful features from buoy data, and channel attention further enhances feature representation. In contrast, other models exhibit performance degradation, suggesting a domain gap between ERA5 and buoy data. Notably, WaveTransNet still performs well among single-channel models, consistent with the ERA5 evaluation results, highlighting the effectiveness of spatial attention for feature extraction from DDM and Aps.
Fig. 11 shows the scatter density plots for channel 1 on the buoy dataset. The models are consistent with those used in Fig. 8. As shown in Fig. 11(g), the fitted line of SCAWaveNet-CD is closest to the ideal line, followed by SCAWaveNet-CI (Fig. 11(f)). WaveTransNet (Fig. 11(c)) and $\text{CNN}_{f}$-CD (Fig. 11(e)) demonstrate similar results: $\text{CNN}_{f}$-CD exhibits a more concentrated distribution with evident underestimation, whereas WaveTransNet shows slightly greater dispersion but less underestimation. $\text{MLP}_{f}$-CD (Fig. 11(d)) and ViT-Wave (Fig. 11(b)) display greater deviations from the ideal line, and suffer from severe underestimation. In summary, SCAWaveNet demonstrates robust performance on the buoy dataset. Although other models exhibit performance degradation, their prediction trends remain consistent with those observed in the ERA5 evaluations.
@@ VI-1 Impact of Input Format
Existing DL-based models lack discussion on input formats during the training process, especially when considering channel information. We investigate two training strategies. First, channel-specific training involves training four separate models using data from each channel independently. Second, channel-general training merges data from all channels to train a single general model. To analyze the impact of these strategies on model performance, we applied both strategies to four representative single-channel models. The results (presented in the Appendix) show that channel-specific models exhibit better performance than channel-general models. Furthermore, channel-general models display performance variations across different channels, suggesting that they cannot learn balanced presentations of each channel. Therefore, we adopt the channel-specific strategy for experimental analysis.
@@ VI-2 The choice of Channel Strategy
This paper introduces three channel strategies. However, only the effects of CD and CI were discussed in the experiments, and the CP strategy was not included. This is because CP involves more diverse and complex modeling approaches. Specifically, CP can be implemented with a fixed number of related channels or a dynamic adaptation mechanism over time, which requires further analysis in future research.
@@ VI-3 Impact of Other Environmental Factors
Beyond wind speed, other environmental factors such as rainfall, wind direction, sea surface temperature, and salinity also affect the model performance. Since this paper focuses on analyzing the impact of channel information on SWH modeling, only the influence of wind speed on SCAWaveNet is considered. Future research can further examine the effects of these factors on different modeling strategies.
@@ VI-4 Data Imbalance
CYGNSS data exhibit a highly imbalanced distribution, with over 80% of samples within the 1-3 m. Consequently, the model primarily learns features from these common cases. For rare samples (e.g., extreme SWH values), it tends to predict values close to the dominant SWH, leading to underestimation under extreme conditions. Although SCAWaveNet performs better than other models on rare samples, this issue still exists. Future work can explore specialized approaches to address this imbalance.
@@ VII Conclusion
Motivated by the data acquisition principles of CYGNSS, this study first applies channel information to the SWH retrieval task. Leveraging attention mechanisms for dynamic feature adaptation and long-range dependency modeling, we propose a spatial-channel attention-based network, namely SCAWaveNet, aimed at better extracting informative features from the data. Specifically, patch embeddings from each channel of DDMs are treated as independent attention heads to enable joint modeling of spatial and channel dependencies. Additionally, a lightweight attention module is applied to APs, which assigns attention weights along both spatial and channel dimensions. The final merged features integrate spatial and channel-level information from DDMs and APs. Unlike single-channel models, SCAWaveNet employs a four-channel architecture, using data from four channels as inputs and predicting SWH for each channel. This design incorporates richer observations and enables the simultaneous retrieval of four SWH values. Experimental results show that SCAWaveNet outperforms state-of-the-art models on four-channel ERA5 and buoy data, demonstrating superior accuracy and robustness. In particular, the model exhibits significant advantages in the 0–1 m range, with regional validation in the Gulf of Mexico, South China Sea, Philippine Sea, and Coral Sea supporting this finding. We also explore the impact of different channel strategies on the four-channel model. The model using the CD strategy performs better than the CI strategy, demonstrating the importance of channel information interaction. Furthermore, we consider the impact of wind speed on model performance. Wind speed improves model accuracy by reducing underestimation at high SWH values.

### v2

@@ V-C NDBC Buoy Data Evaluation
The Buoy data serve as a critical source of ocean observations. Unlike model-generated data, buoy measurements are continuously collected via onboard sensors and instruments, offering higher accuracy and better representation of real-world marine conditions. Therefore, buoy-derived SWH measurements are widely regarded as a reliable reference for ocean state assessment, and it is necessary to evaluate model’s generalization and robustness with buoy data. In this study, NDBC buoy data are utilized for model evaluation. Specifically, 42 NDBC buoys were selected, as illustrated in Fig. 10, where red dots mark their locations, predominantly distributed along the eastern and western coasts of America and near the Hawaiian Islands. These data align with the temporal range of the test set. Data from each channel are individually matched based on close temporal and spatial proximity, and then merged into four-channel samples, yielding a total of 4,438 four-channel data samples (17,752 data points).
@@
Figure 11: Scatter plots of model predictions for CYGNSS-Buoy dataset: (a) CNN-ConvLSTM, (b) ViT-Wave, (c) WaveTransNet, (d) $\text{MLP}_{f}$-CD, (e) $\text{CNN}_{f}$-CD, (f) SCAWaveNet-CI, (g) SCAWaveNet-CD, (h) SCAWaveNet-CD-wind.
TABLE V: Performance comparison of models on CYGNSS-Buoy data
@@
Due to variations in input format across different models, we adopted distinct strategies for model evaluation. For single-channel models, data from all channels were merged for evaluation. For four-channel models, we used four-channel data to test the model and then flattened the prediction by channel to calculate the final results. The evaluation results of all models are shown in Table V. As a global reanalysis dataset, ERA5 exhibits high consistency with the NDBC buoy data, with an average CC of 0.954 across four channels. This result confirms the reliability of using ERA5 as a reference for model evaluation. Among the evaluated results, SCAWaveNet-CD achieves the best performance on buoy data, with an average RMSE of 0.432 m and a CC of 0.754. SCAWaveNet-CI closely follows, with an RMSE of 0.448 m and a CC of 0.741. These findings indicate that the SCAWaveNet architecture effectively captures meaningful features from buoy data, with channel attention further enhancing feature representation. In contrast, other models exhibit performance degradation, suggesting a domain gap between ERA5 and buoy data. Notably, WaveTransNet still performs well among single-channel models, consistent with the ERA5 evaluation results, highlighting the effectiveness of spatial attention for feature extraction from DDM and APs.
Fig. 11 shows the scatter density plots on the CYGNSS-Buoy test set. The models are consistent with those used in Fig. 8. As shown in Fig. 11(g), the fitted line of SCAWaveNet-CD is closest to the ideal line, followed by SCAWaveNet-CI (Fig. 11(f)). WaveTransNet (Fig. 11(c)) and $\text{CNN}_{f}$-CD (Fig. 11(e)) demonstrate similar trends: $\text{CNN}_{f}$-CD exhibits a more concentrated distribution with evident underestimation, whereas WaveTransNet shows slightly greater dispersion but less underestimation. $\text{MLP}_{f}$-CD (Fig. 11(d)) and ViT-Wave (Fig. 11(b)) display greater deviations from the ideal line, and suffer from severe underestimation. In summary, SCAWaveNet demonstrates robust performance on the buoy dataset. Although other models exhibit performance degradation, their prediction trends remain consistent with those observed in the ERA5 evaluations.
@@ VI-1 Impact of Input Format
Existing DL-based models lack discussion of input formats during the training process, especially when considering channel information. We investigate two training strategies. First, channel-specific training involves training four separate models, each using data from a single channel independently. Second, channel-general training merges data from all channels to train a single general model. To analyze the impact of these strategies on model performance, we applied both strategies to four representative single-channel models. The results (presented in the Appendix) show that channel-specific models exhibit better performance than channel-general models. Furthermore, channel-general models display performance variations across different channels, suggesting that they fail to learn balanced presentations for each channel. Therefore, we adopt the channel-specific strategy for experimental analysis.
@@ VI-2 The choice of Channel Strategy
This paper introduces three channel strategies. However, only the effects of CD and CI were discussed in the experiments, while the CP strategy was not included. This is because CP involves more diverse and complex modeling approaches. Specifically, CP can be implemented either with a fixed number of related channels or a dynamic adaptation mechanism over time, which requires further investigation in future research.
@@ VI-3 Impact of Other Environmental Factors
Beyond wind speed, other environmental factors such as rainfall, wind direction, sea surface temperature, and salinity also affect the model performance. Since this paper focuses on analyzing the impact of channel information on SWH modeling, only the influence of wind speed on SCAWaveNet is considered. Future research could further examine the effects of these factors on different modeling strategies.
@@ VI-4 Data Imbalance
CYGNSS data exhibit a highly imbalanced distribution, with over 80% of samples within the 1-3 m. Consequently, the model primarily learns features from these frequent cases. For rare samples (e.g., extreme SWH values), it tends to predict values close to the dominant SWH, leading to underestimation under extreme conditions. Although SCAWaveNet performs better than other models on rare samples, this issue still exists. Future work can explore specialized techniques to address this imbalance.
@@ VII Conclusion
Motivated by the data acquisition principles of CYGNSS, this study first applies channel information to the SWH retrieval task. Leveraging attention mechanisms for dynamic feature adaptation and long-range dependency modeling, we propose a spatial-channel attention-based network, namely SCAWaveNet, aimed at better extracting informative features from the data. Specifically, patch embeddings from each channel of DDMs are treated as independent attention heads to enable joint modeling of spatial and channel dependencies. Additionally, a lightweight attention module is applied to APs, which assigns attention weights along both spatial and channel dimensions. The final merged features integrate spatial and channel-level information from both DDMs and APs. Unlike single-channel models, SCAWaveNet employs a four-channel architecture, using data from all four channels as inputs and predicting SWH for each channel. This design incorporates richer observations and enables the simultaneous retrieval of four SWH values. Experimental results show that SCAWaveNet outperforms state-of-the-art models on four-channel ERA5 and buoy data, achieving superior accuracy and robustness. In particular, the model exhibits significant advantages in the 0–1 m range, with regional validation conducted on the Gulf of Mexico, South China Sea, Philippine Sea, and Coral Sea supporting this finding. We also explore the impact of different channel strategies on the four-channel model. The model using the CD strategy outperforms the CI strategy, highlighting the importance of channel information interaction. Furthermore, we examine the impact of wind speed on model performance. Wind speed improves model accuracy by reducing underestimation at high SWH values.